OpenStudy (anonymous):

I need help solvin the equation x/7-8=-10

OpenStudy (anonymous):

okay I am ready and reporting to you help..

OpenStudy (tiffanymak1996):

x/7 =-10+8

OpenStudy (tiffanymak1996):

x/7 = -2 x=7 times -2 x=-14

OpenStudy (anonymous):

we are solving for x right?

OpenStudy (anonymous):

yes

OpenStudy (tiffanymak1996):

check your answer: x=-14 x/7 -8 = (-14)/7 -8 =-2-8 =-10

OpenStudy (anonymous):

no no you can't do that when x/7 no

OpenStudy (tiffanymak1996):

why?

OpenStudy (anonymous):

whenever you see / it means divide

OpenStudy (anonymous):

so therefore we x divided by 7

OpenStudy (tiffanymak1996):

yeah, i know, when you change sides, it becomes mutiplying

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

???

OpenStudy (tiffanymak1996):

but you have to change sides with addition (or subtraction first before you can change sides with division)

OpenStudy (anonymous):

so it should be 7x=-14

OpenStudy (anonymous):

then divide by 7?

OpenStudy (anonymous):

no it's divided here so that can not happen

OpenStudy (anonymous):

anytime you have something divided by something you times it

OpenStudy (anonymous):

in order for the x or whatever you are solving for.

OpenStudy (anonymous):

ok, let's do this one for clarity.......1-n/5=12

OpenStudy (anonymous):

subtract -1 from both sides...

OpenStudy (anonymous):

you have n/5=11

OpenStudy (anonymous):

so the answer is 55?

OpenStudy (tiffanymak1996):

negative sign!

OpenStudy (tiffanymak1996):

you should have -n/5 =11

OpenStudy (anonymous):

-11?

OpenStudy (anonymous):

oh, I see!!!

OpenStudy (anonymous):

guys if you are doing this right than have X=??

OpenStudy (anonymous):

-55

OpenStudy (tiffanymak1996):

1- n/5 =12 1 -1 -n/5 =12-1 -n/5 =11 -n=55 n=-55

OpenStudy (anonymous):

yipee!

OpenStudy (tiffanymak1996):

gr8! just make sure you don't get confused with the signs

OpenStudy (anonymous):

oh I see what you are doing now, I go step by step that is why I was thrown off by what you did

OpenStudy (tiffanymak1996):

sry!

OpenStudy (anonymous):

6t +3=-7

OpenStudy (anonymous):

subtract 3 from both sides

OpenStudy (tiffanymak1996):

and you can always check your answers by substituting x=? in you equation

OpenStudy (anonymous):

6t=-10

OpenStudy (anonymous):

divide by 6?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

answer is -2?

OpenStudy (tiffanymak1996):

what is -10 divided by 6?

OpenStudy (anonymous):

something isn't right, that is not one of my answwer choices

OpenStudy (anonymous):

10 goes into 2 five and 2 goes into 6 3 so the answer is -5/3

OpenStudy (tiffanymak1996):

well your steps are all correct...

OpenStudy (anonymous):

I didn't have the lowest denominator

OpenStudy (tiffanymak1996):

usually the answers are in simplest form, so simplify your fractions

OpenStudy (tiffanymak1996):

what choices do you have?

OpenStudy (anonymous):

Tiffany I am so sorry about this, I was trying to walk through the steps here so you could ask questions. if you have this I will go

OpenStudy (anonymous):

I have -5/3 as an answer

OpenStudy (anonymous):

no before you said -2

OpenStudy (anonymous):

right, but godorovg came up with -5/3 and I have that as an answer choice

OpenStudy (anonymous):

I did not simplify

OpenStudy (anonymous):

unless it says to do it I would leave as is

OpenStudy (anonymous):

15=-15-8u

OpenStudy (anonymous):

add 15 to both sides

OpenStudy (anonymous):

30-8u

OpenStudy (anonymous):

30=-8u

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

divide by -8?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

15/-4?

OpenStudy (anonymous):

-15/4?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

I need help with this one 0=1/6y+8=

OpenStudy (anonymous):

subtract 8 from both sides

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

0-8 is 0?

OpenStudy (anonymous):

-8?

OpenStudy (tiffanymak1996):

yes, -8

OpenStudy (anonymous):

-8=1/6y

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

can I help here, this problem you are making it harder than it has be

OpenStudy (tiffanymak1996):

because the variable is in the denominator, therefore we have to switch it to the other side

OpenStudy (tiffanymak1996):

-8=1/6y -8(6y)=1/6y times 6y

OpenStudy (tiffanymak1996):

so you multiply both sides with 6y

OpenStudy (anonymous):

yes that is right, although when you first start this collect all a like terms and move then to one side and than times through the number being divided

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

okay let us take this one x/7-8=-10 now here is how I would do this First take the -8 and add 8 to the -10 so now we have this x/7=-10+8 than take this and do the sub to get -2 so now we have x/7=-2 once you have than take the seven ant times to the two so now we have X= -14 like that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-8=1/6y?

OpenStudy (anonymous):

okay now you tell me

OpenStudy (anonymous):

multiply both sides by 1/6?

OpenStudy (tiffanymak1996):

you multiply both sides with 6y

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it is 1/6 not just six

OpenStudy (anonymous):

so y=1/6x8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or 48?

OpenStudy (tiffanymak1996):

it should be y=1/-48

OpenStudy (anonymous):

oh yes, because it was a -8

OpenStudy (tiffanymak1996):

since: 1/6y =-8 1/(6 times-8) =y so, y=1/-48

OpenStudy (anonymous):

yes it that is the y=1/-48

OpenStudy (anonymous):

okay, thanks guys! I'm tired...going to bed, goodnight! Thanks for your help!

OpenStudy (anonymous):

hey no problem glad I could you help you :)

OpenStudy (tiffanymak1996):

gd nite!

OpenStudy (anonymous):

tiffany great job!!

OpenStudy (tiffanymak1996):

u too!